OpenStudy (aravindg):

4^x =65536 how do i get x manually?

OpenStudy (anonymous):

Do you what is \( 2^{16} \)?

OpenStudy (lgbasallote):

one way is to use logarithms...another way is to expres 65536 in power 4 like 4^8

OpenStudy (callisto):

4^x =65536 log both sides xlog4 = log65536 x = 8

OpenStudy (anonymous):

know*

OpenStudy (mayliz):

x=8 4^8=65536

OpenStudy (aravindg):

hey did we divide both the logs in last step??@Callisto

OpenStudy (amistre64):

manually, youd use your finger to type the keys on the calculator :)

OpenStudy (aravindg):

i cant guess its 4^8 lol

OpenStudy (aravindg):

@amistre calculator is autonatic! i said "manually"

sam (.sam.):

@amistre64 that's what i do :)

OpenStudy (aravindg):

hey so hw do we get it using power of 2??@FoolForMath

OpenStudy (aravindg):

oh (2^2)^x !!!

OpenStudy (aravindg):

2^2x

OpenStudy (aravindg):

hmm... well that log part is new information to me.i think its intersting way

OpenStudy (aravindg):

take log on bothg sides and get the result!!

OpenStudy (aravindg):

hey i frgt to ask what is the base???

OpenStudy (callisto):

2|65536 2|32768 2|16384 2|8192 2|4096 2|2048 2|1024 2|512 2|256 2|128 2|64 2|32 2|16 2|8 2|4 2|2 1 count the no of 2 and you'll get the power with base 2 and if this is what you mean by manually

OpenStudy (callisto):

it a common log

OpenStudy (amistre64):

\[\int \frac{1}{1+x}dx = ln(1+x)\] 1-x+x^2-x^3+x^4-x^5 .... -------------------- 1+x ) 1 (1+x) ------- -x (-x-x^2) --------- +x^2 \[\int (1-x+x^2-x^3+x^4...)dx = ln(1+x) =x-\frac{x^2}{2}+\frac{x^3}{3}-\frac{x^4}{4}+\frac{x^5}{5}...\]

OpenStudy (aravindg):

common log means base e??

OpenStudy (callisto):

base 10

OpenStudy (callisto):

i think e is that natural log

OpenStudy (amistre64):

\[4^x =65536\] \[ln(4^x =65536)\] \[ln(4^x) =ln(65536)\] \[xln(4) =ln(65536)\] \[x =\frac{ln(65536)}{ln(4)}\]

OpenStudy (amistre64):

use the poly of ln(1+x) to manually determing the rest ;)

OpenStudy (aravindg):

so natural log wont work??

OpenStudy (callisto):

it works also

OpenStudy (anonymous):

all logs are the same, you can use any one

OpenStudy (amistre64):

all logs lead to the bayou

OpenStudy (anonymous):

how's bayou?